# EFUF: Efficient Fine-Grained Unlearning Framework for Mitigating Hallucinations in Multimodal Large Language Models

### v2

@@ Abstract
Multimodal large language models (MLLMs) have attracted increasing attention in the past few years, but they may still generate descriptions that include objects not present in the corresponding images, a phenomenon known as object hallucination. To eliminate hallucinations, existing methods manually annotate paired responses with and without hallucinations, and then employ various alignment algorithms to improve the alignment capability between images and text. However, they not only demand considerable computation resources during the finetuning stage but also require expensive human annotation to construct paired data needed by the alignment algorithms. To address these issues, we propose an efficient fine-grained unlearning framework (EFUF), which performs gradient ascent utilizing three tailored losses to eliminate hallucinations without paired data. Extensive experiments show that our method consistently reduces hallucinations while preserving the generation quality with modest computational overhead. Our code and datasets will be publicly available.
@@ 1 Introduction
Recent methods for mitigating multimodal hallucination can be divided into two categories: revision-based methods (Lee et al., 2023; Zhou et al., 2023; Yin et al., 2023; Wang et al., 2023) and finetuning-based methods (Sun et al., 2023; Yu et al., 2023; Liu et al., 2023a; Zhao et al., 2023; Jiang et al., 2023). Revision-based strategies correct generated content through external expert review or self-reflection during inference stage. However, they usually require additional inference steps with increased costs and delay (Yu et al., 2023). Furthermore, each task demands specific procedure and prompt (Xu et al., 2024), adding to the complexity of implementation. Overcoming these drawbacks, finetuning-based approaches are proposed to adjust the model directly through specialized datasets and preference alignment algorithms. These algorithms, including RLHF (Sun et al., 2023; Liu et al., 2023a), DPO (Yu et al., 2023; Zhao et al., 2023; zhou2024aligning) and contrastive learning (Jiang et al., 2023), enhance the congruence between text and images, leading to improved alignment. Although they have achieved good performance, two critical issues emerge:
@@
The key to applying the unlearning algorithm is how to curate positive and negative samples, i.e., distinguish between real and hallucinated objects, in a manner that is both cost-effective and reliable. Intuitively, the similarity between objects and their corresponding images can act as an indicator for hallucinations, since the image contains real objects but not the hallucinated ones. In this work, we propose to utilize the CLIP model (Radford et al., 2021) to evaluate text-image congruence. Trained on a vast corpus of text-image pairs, CLIP stands as a robust tool to help identify hallucinations.
@@ 2.1 Hallucination Mitigation for MLLM
(1) Revision-based methods. They employ external experts or self-reflection strategies to correct generated content during the inference phase, thereby reducing hallucinations. For example, LURE (Zhou et al., 2023) utilizes manually-crafted features to detect hallucinations and therefore revises the generated text. Woodpecker (Yin et al., 2023) proposes to post-edit hallucinations by combining the output of MLLMs and a more accurate expert VQA model using GPT-3.5. VIGC (Wang et al., 2023) iteratively refines the instruction data using generation and correction framework. VOLCANO (Lee et al., 2023) trains the MLLM to give self-feedback, and then performs self-reflection on the original generated text according to the feedback. Although these methods do not need to train the model, they require additional inference steps with increased costs and delay (Yu et al., 2023), and specific procedure and prompt must be designed for each task (Xu et al., 2024).
(2) Finetuning-based methods. Overcoming the potential drawbacks of the first category, these methods involve crafting specific datasets and finetuning the model, aiming for better alignment between images and text. For instance, LLaVA-RLHF (Sun et al., 2023) first adopts RLHF to mitigate hallucinations. Based on this work, RLHF-V (Yu et al., 2023) introduces fine-grained alignment by manually correcting the outputs of MLLMs. Beyond standard RLHF, some works utilize other improved algorithms for better efficiency, e.g., DPO (Zhao et al., 2023; zhou2024aligning), instruction tuning (Liu et al., 2023a), and contrastive learning (Jiang et al., 2023). However, these methods require expensive manually annotated paired data, and most of them also demand substantial computational resources during the finetuning stage. Therefore, in this work, we focus on reducing the data and computation requirements.
@@ 6 Results and Analysis
Table 4: Performance comparison of different hallucination mitigation methods on LLaVA. Best scores are in bold and second bests are underlined.
@@ 6.2 Ablation Study
6.3 Comparison with Other Hallucination Mitigation Methods
To further evaluate the performance of EFUF, we compare it with other methods tailored to hallucination mitigation. These include LLaVA-RLHF (Sun et al., 2023), HA-DPO (Zhao et al., 2023), and POVID (zhou2024aligning), which are all evaluated using their officially released checkpoints. We benchmark EFUF against these methods on the LLaVA model, since their checkpoints are all based on LLaVA. The results are shown in Table 4.
Compared to other hallucination mitigation methods, EFUF demonstrates comparable or superior performance, while requiring minimal data construction cost and training resources among all. Additionally, our improvements in generation quality are on par with RLHF-based methods, which typically demand expensive human annotations and significant computations. These outcomes highlight our method’s effectiveness and efficiency.
6.4 More Analyses
To underscore the effectiveness of our approach, additional analyses are provided in the appendix. A comparison of the training expenses for EFUF and alternative finetuning-based methods is given in Appendix B. This comparison highlights that EFUF reduces the resources needed for both dataset construction and training process by a large margin. Further experimental evidence, presented in Appendix C, illustrates that EFUF complements and enhances the performance of existing hallucination mitigation strategies. We also explore the impact of varying weights as hyper-parameters in Appendix D. Finally, a case study detailed in Appendix E quantitatively evaluates the generated text under different methods, showcasing the distinct advantages of our proposed solution.
@@ A.3.2 Metrics on Generation Quality
Appendix B Training Cost
Appendix C EFUF is beneficial to other hallucination mitigation methods
@@
Appendix D Effects of different weight
@@
Appendix E Case Study

### v3

@@ Abstract
Multimodal large language models (MLLMs) have attracted increasing attention in the past few years, but they may still generate descriptions that include objects not present in the corresponding images, a phenomenon known as object hallucination. To eliminate hallucinations, existing methods manually annotate paired responses with and without hallucinations, and then employ various alignment algorithms to improve the alignment capability between images and text. However, they not only demand considerable computation resources during the finetuning stage but also require expensive human annotation to construct paired data needed by the alignment algorithms. To address these issues, we propose an efficient fine-grained unlearning framework (EFUF), which performs gradient ascent utilizing three tailored losses to eliminate hallucinations without paired data. Extensive experiments show that our method consistently reduces hallucinations while preserving the generation quality with modest computational overhead. Our code and datasets are available at https://github.com/starreeze/efuf.
@@ 1 Introduction
Recent methods for mitigating multimodal hallucination can be divided into two categories: inference-based methods (Lee et al., 2023; Zhou et al., 2023; Yin et al., 2023; Wang et al., 2023; Sicong Leng, 2023; Wang et al., 2024; Chen et al., 2024) and finetuning-based methods (Sun et al., 2023; Yu et al., 2023; Liu et al., 2023a; Zhao et al., 2023; Jiang et al., 2023). Inference-based methods correct or restrict generated content through external expert review, self-reflection or decoding strategies during inference stage. However, they usually require additional inference steps with increased costs and delay (Yu et al., 2023). Furthermore, each task demands specific procedure or prompt (Xu et al., 2024), adding to the complexity of implementation. Overcoming these drawbacks, finetuning-based approaches are proposed to adjust the model directly through specialized datasets and preference alignment algorithms. These algorithms, including RLHF (Sun et al., 2023; Liu et al., 2023a), DPO (Yu et al., 2023; Zhao et al., 2023; Zhou et al., 2024) and contrastive learning (Jiang et al., 2023), enhance the congruence between text and images, leading to improved alignment. Although they have achieved good performance, two critical issues emerge:
@@
The key to applying the unlearning algorithm is how to curate positive and negative samples, i.e., distinguish between real and hallucinated objects, in a manner that is both cost-effective and reliable. Intuitively, the similarity between objects and their corresponding images can act as an indicator for hallucinations, since the image contains real objects but not the hallucinated ones. Inspired by Zhao et al. (2024), we propose to utilize the CLIP model (Radford et al., 2021) to evaluate text-image congruence. Trained on a vast corpus of text-image pairs, CLIP stands as a robust tool to help identify hallucinations.
@@ 2.1 Hallucination Mitigation for MLLM
(1) Inference-based methods. They employ external experts, self-reflection framework or decoding strategies to constrain or modify generated content during the inference phase, thereby reducing hallucinations. For example, LURE (Zhou et al., 2023) utilizes manually-crafted features to detect hallucinations and therefore revises the generated text. Woodpecker (Yin et al., 2023) proposes to post-edit hallucinations by combining the output of MLLMs and a more accurate expert VQA model using GPT-3.5. VIGC (Wang et al., 2023) iteratively refines the instruction data using generation and correction framework. VOLCANO (Lee et al., 2023) trains the MLLM to give self-feedback, and then performs self-reflection on the original generated text according to the feedback. VCD (Sicong Leng, 2023) first introduces contrastive decoding in MLLM by disturbing the visual inputs and calculate visual uncertainty to restrict the generation of hallucinated tokens. ICD (Wang et al., 2024) utilizes disturbance on instructions instead of images. HIO (Chen et al., 2024) employs a hallucinated model to further widen the gap between hallucinated and correct tokens, achieving better contrastive outcomes. Although these methods do not need to train the model, they require additional inference steps with increased costs and delay (Yu et al., 2023), and specific procedure and prompt must be designed for each task (Xu et al., 2024).
(2) Finetuning-based methods. Overcoming the potential drawbacks of the first category, these methods involve crafting specific datasets and finetuning the model, aiming for better alignment between images and text. For instance, LLaVA-RLHF (Sun et al., 2023) first adopts RLHF to mitigate hallucinations. Based on this work, RLHF-V (Yu et al., 2023) introduces fine-grained alignment by manually correcting the outputs of MLLMs. Beyond standard RLHF, some works utilize other improved algorithms for better efficiency, e.g., DPO (Zhao et al., 2023; Zhou et al., 2024), instruction tuning (Liu et al., 2023a), and contrastive learning (Jiang et al., 2023). However, these methods require expensive manually annotated paired data, and most of them also demand substantial computational resources during the finetuning stage. Therefore, in this work, we focus on reducing the data and computation requirements.
@@ 6 Results and Analysis
Table 4: Performance comparison of different hallucination mitigation methods for LLaVA on metrics measuring hallucination rate and generation quality. Best scores are in bold and second bests are underlined.
Method MME GQA SQA QBench
LLaVA 1491 63.0 66.9 59.2
   + RLHF 1212 48.4 65.4 53.0
   + HADPO 1441 61.2 67.2 58.6
   + POVID 1438 61.9 68.4 59.2
   + EFUF 1468 63.2 66.4 59.3
Table 5: Performance comparison of different hallucination mitigation methods for LLaVA on metrics measuring VQA and reasoning capability.
@@ 6.2 Ablation Study
6.3 Comparison with Other Methods
To further evaluate the performance of EFUF, we compare it with other methods tailored to hallucination mitigation. These include LLaVA-RLHF (Sun et al., 2023), HA-DPO (Zhao et al., 2023), and POVID (Zhou et al., 2024), which are all evaluated using their officially released checkpoints. We benchmark EFUF against these methods on the LLaVA model, since their checkpoints are all based on LLaVA.
Hallucination Rate & Generation Quality.
We measure EFUF’s generation quality along with hallucination rate in Table 4. Compared to other hallucination mitigation methods, EFUF demonstrates comparable or superior performance, while requiring minimal data construction cost and training resources among all. Additionally, our improvements in generation quality are on par with RLHF-based methods, which typically demand expensive human annotations and significant computations. These outcomes highlight our method’s effectiveness and efficiency.
VQA & Reasoning Capability.
To provide a more holistic evaluation of EFUF, we also assessed its performance on VQA and reasoning tasks. We employed benchmarks such as MME (Fu et al., 2024), GQA (Hudson and Manning, 2019), ScienceQA (Lu et al., 2022), and QBench (Wu et al., 2024). Table 5 reports the results for the baseline model, EFUF, and competing methods. EFUF demonstrates modest performance fluctuation across these benchmarks compared to other hallucination mitigation strategies, indicating that our method does not negatively affect VQA and reasoning capabilities.
6.4 Training Cost
6.5 Additional Analyses
To further substantiate the effectiveness of EFUF, we provide extensive supplementary analyses in the appendices. As presented in Appendix B, EFUF complements and enhances the performance of existing hallucination mitigation strategies. We also explore the impact of varying weights as hyper-parameters in Appendix C. Finally, a case study detailed in Appendix D quantitatively evaluates the generated text under different methods, showcasing the distinct advantages of our proposed solution.
@@ 7 Conclusion
Acknowledgements
We would like to thank the anonymous reviewers for their constructive comments. This work was supported by the National Natural Science Foundation of China (No. 62206126 and No. 61976114).
@@ A.3.2 Metrics on Generation Quality
Appendix B EFUF is beneficial to other hallucination mitigation methods
@@
Appendix C Effects of different weight
@@
Appendix D Case Study